# ContactHandover: Contact-Guided Robot-to-Human Object Handover

### v1

@@ I INTRODUCTION
Object handover is a key step towards natural human and robot collaboration [8, 27]. Robot-to-human handover, in particular, has wide applications in a lot of practice scenarios, from handing tools to workers in factory to fetching daily objects for elders at home. A handover process typically involves a grasping phase in which the robot picks up the target object, and a delivery phase where the robot approaches the human and moves the object to a pose that is accessible and ergonomic for the human receiver to grasp and use in subsequent tasks.
@@
To address these challenges, we propose a robot-to-human handover system, ContactHandover, which uses 3D contact maps to model diverse human preferences when receiving objects. Our system contains two phases. In the object grasping phase, the system predicts 6-DoF robot grasp poses and a human contact map for the object, re-ranks the grasp poses to penalize those that occlude human contacts, and executes the highest ranking grasp. During the delivery phase, we compute the robot end-effector position and orientation that both minimizes the human arm joint torques and displacements when receiving the object, and maximizes human contact points that are close to the human.
To evaluate our result quantitatively, we propose two computational metrics that align with previous work’s discovery on ergonomic object delivery poses for human receivers [28][3]. The visibility metric measures the percentage of human contact points from the receiver’s viewpoint. The reachability metric measures the percentage of human contact points reachable from the receiver without obstruction from the robot’s embodiment.
Figure 1: Contact-Guided Robot to Human Object Handover. We propose a robot-to-human handover system with two phases: (a) contact-guided grasping and (b) object delivery. (a) During grasping, the robot predicts 6-DoF grasp pose candidates and human contact points (denoted in red) for the object on the table, and selects a grasp pose that maximizes stability while minimizing occlusion of human contact points. (b) During delivery, the robot computes a handover location and orientation that minimizes human arm joint torque and displacements and maximizes human contact points close to the receiver.
@@
  • A contact-guided grasp selection algorithm that considers both grasp stability of the robot and contact preferences of the human receiver.
  • An object delivery algorithm that computes a robot end effector pose that minimizes human arm joint torques as well as displacements, and maximizes contact points close to the human receiver.
  • Two benchmark metrics (visibility, reachability) that quantitatively evaluates the quality of a handover pose.
@@ II RELATED WORK
Grasp Pose Prediction has been a long standing task in robotics. Early works use analytical methods to plan and execute stable grasps, which are limited in real world applications given the assumption of known object geometry [13, 24, 14]. More recently, data-driven approaches use convolutional neural networks to learn grasp affordances directly from top-down RGB-D images [21, 34, 15], or use generative models such as VAE to sample grasp candidates from object point cloud and then filter based on the grasping quality [25, 30]. In addition to predicting stable grasps, other works have studied predicting functional grasps, namely grasping on functional parts of the object to use [6, 35]. During a robot-to-human handover, we identify functional grasps as those that avoid human preferred contact points, and propose a contact-guided grasp ranking method to predict stable grasp that also maximizes available human contact points on the object.
@@
Learning Human Grasp Affordances. During a robot-to-human handover, it is important to generate grasp pose and delivery pose that accommodate human grasp affordances. To model human grasp affordances, one line of works directly learn 3D affordance maps on objects with respect to different intents (e.g. use, handoff) [5, 4]. In particular, ContactDB [5] is a dataset of human contact maps for household objects collected with a thermal camera. Other works model human grasp affordance by predicting hand shape and pose when grasping  [10, 16, 31, 32, 12]. In this work, we use 3D human contact affordance maps as a proxy for human preferences while receiving objects and show that the contact maps can be used to effectively guide robot grasp pose and delivery pose selection during a robot-to-human handover.
@@
Robot-to-Human Object Handover. The main objective of robot-to-human object handover is to deliver the object in a way that maximizes the user’s ease to grasp and convenience to use the object for a subsequent task [26]. Prior works have attempted to predict and maximize human grasp affordance during handover, but they either manually select human grasping part on an object [2], make assumptions about object geometry and hand-design object categories [9, 3, 17], or train on synthetic data that fail to model the complexity of human contacts [19]. Other works use simplifying heuristics such as assuming that the robot grasp is on the opposite side of a predicted human hand pose [23]. A key limitation of this heuristic is that it does not account for scenarios where both the human and robot have to grasp on same side of an object (e.g. knife handle). We propose a novel approach that predicts human contact points on an object with no assumption of object category and use human contact points to guide robot grasp pose and delivery pose selection.
@@
Figure 2: Contact-Guided Grasp Selection. (a)§III-A1 The robot takes RGB-D Images from 16 views around the table and construct a 64*64*64 voxel representation of the object via TSDF fusion; the occupancy grid is then fed into a trained 3D VoxNet to predict human contact maps. (b)§III-A2 The robot takes a partial point cloud observation as input to the pre-trained Contact-GraspNet model to generate a set of 6-DoF robot grasps. (c)§III-A3 The robot executes the grasp with highest score as computed by Equation 2.
@@ III METHOD
Our robot-to-human handover system contains two phases: a contact-guided grasping phase and an object delivery phase. During the grasping phase, given RGB-D observations of an object on table, the system predicts both 6-DoF robot grasp pose candidates and human contact points on the object, and then re-ranks the grasp poses by penalizing poses based on number of human contact points occluded by the robot end effector. During the delivery phase, the system computes a handover position and orientation that minimizes the human arm joint torques and displacements, and minimizes the total distances from contact points to the receiver’s eyes. We will discuss each module in detail below.
@@ III-A Contact-Guided Grasp Selection
A successful handover requires the robot to select a stable grasp location on the object that leaves room for the human receiver’s grasp later. As illustrated in Fig. 2, given RGB-D observations of an object on the table, the robot predicts a set of grasp pose candidates together with a human contact map on the object. A final score is computed for each robot grasp pose candidate by combining the contact confidence of the grasp and percentage of occluded human contact points. The grasp with the highest score is chosen and executed.
@@ III-A2 Robot Grasp Prediction
Contact-GraspNet is a PointNet++ based U-shaped model that takes in a partial point cloud observation of a scene, and for each point $i$, predicts whether it is contacted by the robot gripper during grasping with a confidence score $S(i)$. For each point $i$, the model predicts the 3-DoF grasp orientation and grasp width $w\in R$ of a parallel-yaw gripper. Each robot contact point $i$, together with its 4-DoF grasp representation, can be translated to a 6-DoF robot gripper pose $g$ for grasping. Following [30], we select grasps with confidence $S(g)\geq 0.23$ as robot grasp candidates. We use the predicted confidence score for each grasp $S(g)$ in the re-ranking phase.
@@ III-A3 Robot Grasp Re-Ranking
Clustering Human Contacts: For some objects, like the binoculars, the human use mode typically involves both hands. However, during handovers, a human will usually receive these objects with one hand. Moreover, avoiding all possible contact points might leave few valid robot grasp candidates after re-ranking, leading to grasp failures. To resolve this, we cluster predicted human contact points using DBSCAN[11] based on point spatial density, without no assumptions on the number of clusters. We then choose the largest human contact cluster to compute human contact occlusion and handover orientation in §III-C.
Grasp Re-Ranking: For each grasp $g$, we calculate the percentage of human contact points that are occluded by the robot gripper, denoted as $O(g)$. For each human contact point, we ray trace from the contact point along it’s surface normal and check if the ray collides with the robot gripper at grasp pose $g$. If there’s a collision, then we consider the human contact point to be blocked by the gripper at pose $g$.
$O(g)=\frac{\Big{|}\big{\{}i\in C_{pred}\mid i\textrm{ blocked by }g\big{\}}\Big{|}}{\big{|}C_{pred}\big{|}}$ (1)
@@
Lastly, We re-rank the robot grasps to account for the number of human contacts they occlude. We calculate a final contact score $C(g)$ for each grasp that combines the grasping contact confidence $S(g)$ and the occluded human contacts $O(g)$.
$C(g)=\lambda S(g)-(1-\lambda)O(g)$ (2)
where $\lambda$ controls the weight between grasp confidence and human contact occlusion. We use $\lambda=0.5$ in our experiments. Finally, the robot executes the grasp with the highest score $C(g)$.
@@ III-B Handover Position
During the delivery phase, the robot should hand the object to a point in front of the human that is both reachable and comfortable for the human arm, with respect to the human’s height and pose. For instance, the handover position for a human that is standing should differ from a human that is sitting down. Following previous works [18, 29, 20], ContactHandover computes the point of handover by minimizing human arm joint torques and joint displacement. In addition, we enforce that the handover position is within the vertical range of the human’s torso (below the shoulder and above the waist).
@@
Finally, we can define the total cost of a candidate point as
@@
where $\alpha$ is a hyperparameter to control the weight of the two cost functions.
@@ III-C Handover Orientation
In the delivery stage, the robot must present the object in an orientation that maximizes contacts points close to the human. Specifically, we uniformly sample in the spherical space with a 45 degree granularity and select the object orientations that are kinematically feasible [1] for the robot arm at the computed handover point. For each candidate orientations, ContactHandover computes the total distance between predicted contact points and the human eyes. We estimate the human eye to be located in a position that is exactly 1/2 the height of the human head. The orientation that yields minimum distance is selected.
Note that as in §III-A3, we compute the distance for the largest human contact cluster. For objects with clustered human contacts (e.g. binoculars), we found that orienting one cluster of human contact towards the human results in more natural handovers, since human tends to receive these objects with one hand instead of two hands. We show the qualitative result for the clustering in §V and Figure 5.
@@ IV-A Handover Metrics
We evaluate our system on 27 daily objects from the ContactDB dataset, which consists of 50 contact maps collected from 50 different users for each object. We introduce two quantitative metrics based on the contact maps to evaluate the robot to human handover result.
Metric 1: Human Contact Visibility. To ensure that humans can easily grasp their preferred contact areas upon receiving an object, it is crucial that these areas are visible to the receiver. In particular, the areas where humans prefer to make contact should not be occluded by the robot’s embodiment or by the object itself.
@@
where $CM:i\mapsto c\in\{0,1\}$ indicates if each voxel $i$ on the object surface is touched during a human grasp. $\mathit{P}$ is the set of all voxels in the contact map. $\mathit{V}$ is the set of voxels that are visible from the human’s view without occlusion from the robot embodiment or the object itself. We compute visibility by setting a RGB-D camera at the human receiver’s eye position looking at the object; and any contact points captured in the image are considered visible.
@@
where $\mathit{R}=\{(i:d_{1}(i)<l_{\mathbf{arm}})\cap(d_{2}(i)<d_{2}(\mathbf{gripper}))\}$. $d_{1}$ represents the distance from point $i$ to human shoulder and $l_{\mathbf{arm}}$ is the human’s arm length. $d_{2}(i)$ represents the shortest horizontal distance from point $i$ to the human; $d_{2}(\mathbf{gripper})$ represents the shortest distance between the robot gripper and human. If $d_{2}(i)<d_{2}(\mathbf{gripper})$, then the point $i$ is located between the human and the robot gripper. Higher reachability score indicates more human preferred contact areas are facing towards the human.
Success Rates. For each object, we compute the visibility and reachability score based on the 50 corresponding contact maps and select the median as the final score. We consider a handover to be successful if both visibility and reachability score is above a threshold $k$, where we choose $k=0.5$ for our evaluation.
@@ IV-B Experiment Setup
We set up a revolving RGB-D camera that captures 16 distinct angles around the tabletop to reconstruct the object point cloud and voxel grids; and we simulate a RGB-D camera on the human eye level to observe the handover object and calculate visibility metric.
For each object, the robot first grasps the object on table, then turn around and move to 1.2 meters in front of the human, and finally handover the object in the computed pose. We run the experiment for 27 selected objects from the ContactDB “use” dataset under 5 random seeds and report the average.
@@ IV-C Baselines
Ablation 1 (A1): No Grasp Re-ranking (GR). The robot does not predict and consider human contacts when grasping the object. Instead, the robot only predicts and executes the grasp with the highest confidence score.
@@
Ablation 3 (A4). No optimization. The robot executes the grasp with highest confidence score, approaches and stops in front of the human at the same distance as ContactHandover, while maintaining the same end effector pose after grasping.
Figure 4: Generalize to Unseen Objects. We show ContactHandover’s performance on unseen YCB objects. We show both human contact predictions and handover results on these objects. ContactHandover is able to generalize to (a) objects with unseen shapes (e.g. hammer and knife) and (b) objects with unseen types (e.g. spoon, flat screwdriver, fork and phillips screwdriver). It predicts reasonable human contacts (denoted in red points) around the handles of the objects, picks up and delivers the objects to human with respect to the predicted human contacts
@@
Figure 5: Clustering bimodal human contacts. For objects with bimodal human contact distributions, ContactHandover clusters the contact points and only optimizes human contact points in one cluster. (a) shows the result for ContactHandover, where the robot leaves out and orients one side of the binoculars towards the human. In (b), without clustering, the robot points both side towards the human, resulting in a pose where neither human contact clusters is close to the human.

### v2

@@ I INTRODUCTION
Object handover is a key step towards natural human and robot collaboration [7, 3, 35, 9, 34]. Robot-to-human handover, in particular, has wide applications in a lot of practice scenarios, from handing tools to workers in factory to fetching daily objects for elders at home. A handover process typically involves a grasping phase in which the robot picks up the target object, and a delivery phase where the robot approaches the human and moves the object to a pose that is accessible and ergonomic for the human receiver to grasp and use in subsequent tasks.
@@
To address these challenges, we propose a robot-to-human handover system, ContactHandover, which uses 3D contact maps to model diverse human preferences when receiving objects. Our system contains two phases. In the object grasping phase, the system predicts 6-DoF robot grasp poses and a human contact map for the object, re-ranks the grasp poses to penalize those that occlude human contacts, and executes the highest ranking grasp. During the delivery phase, we compute the robot end-effector position and orientation that both minimizes the human arm joint torques and displacements when receiving the object, as well as the distances from contact points to the receiver eyes’ location.
To evaluate our result quantitatively, we propose two computational metrics, visibility and reachability, that align with previous work’s discovery on ergonomic object delivery poses for human receivers [35, 3]. The visibility metric measures the percentage of human contact points from the receiver’s viewpoint. The reachability metric measures the percentage of human contact points reachable from the receiver without obstruction from the robot’s embodiment. Finally, a handover is considered ”successful” if both the visibility and reachability metric exceed a threshold.
Figure 1: Contact-Guided Robot to Human Object Handover. We propose a robot-to-human handover system with two phases: (a) contact-guided grasping and (b) object delivery. (a) During grasping, the robot predicts 6-DoF grasp poses and human contact points (denoted in red) for the object, and selects a grasp pose that maximizes stability while minimizing contact points occlusions. (b) During delivery, the robot computes a handover location and orientation that minimizes human arm joint torque and displacements, as well as the distance between contact points and the human.
@@
  • A contact-guided grasp selection algorithm that accounts for both grasp stability of the robot and contact preferences of the human receiver.
  • An object delivery algorithm that computes the robot end effector pose by considering the human’s arm comfort when receiving the object and minimizing the distance between the contact points and the human.
  • Two benchmark metrics (visibility, reachability) that quantitatively evaluate a handover pose.
@@ II RELATED WORK
Grasp Pose Prediction has been a long standing task in robotics. Early works use analytical methods to plan and execute stable grasps, which are limited in real world applications given the assumption of known object geometry [16, 31, 18]. More recently, data-driven approaches use convolutional neural networks to learn grasp affordances directly from top-down RGB-D images [28, 41, 19], or use generative models such as VAE to sample grasp candidates from object point cloud and then filter based on the grasping quality [32, 37]. In addition to predicting stable grasps, other works have studied predicting functional grasps, namely grasping on functional parts of the object to use [6, 42, 17, 43, 27]. In the robot-to-human handover context, we identify functional grasps as those that avoid human preferred contact points, and propose a contact-guided grasp ranking method to predict stable grasp that also maximizes available human contact points on the object.
@@
Learning Human Grasp Affordances. During a robot-to-human handover, it is important to generate grasp pose and delivery pose that accommodate human grasp affordances. To model human grasp affordances, one line of works directly learn 3D affordance maps on objects with respect to different intents (e.g. use, handoff) [5, 4, 13]. In particular, ContactDB [5] is a dataset of human contact maps for household objects collected with a thermal camera. Other works model human grasp affordance by predicting hand shape and pose when grasping  [11, 20, 38, 39, 15]. In this work, we use 3D human contact affordance maps as a proxy for human preferences while receiving objects and show that the contact maps can be used to effectively guide robot grasp pose and delivery pose selection during a robot-to-human handover.
@@
Robot-to-Human Object Handover. The main objective of robot-to-human object handover is to deliver the object in a way that maximizes the user’s ease to grasp and convenience to use the object for a subsequent task [33]. Prior works have attempted to predict and maximize human grasp affordance during handover, but they either manually select human grasping part on an object [2], make assumptions about object geometry and hand-design object categories [10, 3, 21], or train on synthetic data that fail to model the complexity of human contacts [23]. Other works use simplifying heuristics such as assuming that the robot grasp is on the opposite side of a predicted human hand pose [30]. But this heuristic does not account for scenarios where both the human and robot have to grasp on same side of an object (e.g. knife handle). We propose a novel approach that predicts human contact points on an object with no assumption of object category and use human contact points to guide robot grasp pose and delivery pose selection.
@@
Figure 2: Contact-Guided Grasp Selection. (a)§III-A1 The robot takes RGB-D Images from 16 views around the table and construct a $64^{3}$ voxel representation of the object via TSDF fusion; the occupancy grid is then fed into a trained 3D VoxNet to predict human contact maps. (b)§III-A2 The robot takes a partial point cloud observation as input to the pre-trained Contact-GraspNet model to generate a set of 6-DoF robot grasps. (c)§III-A3 The robot executes the grasp with highest score as computed by Equation 2.
@@ III METHOD
ContactHandover contains two phases: a contact-guided grasping phase and an object delivery phase. During the grasping phase, given RGB-D observations of an object on table, the system predicts both 6-DoF robot grasp pose candidates and human contact points on the object, and then re-ranks the grasp poses by penalizing grasps poses based on number of human contact points occluded by the robot end effector. During the delivery phase, the system computes a handover position and orientation that minimizes the human arm joint torques and displacements, and minimizes the total distances from contact points to the receiver’s eyes location. We will discuss each module in detail below.
@@ III-A Contact-Guided Grasp Selection
A successful handover requires the robot to select a stable grasp location on the object that leaves room for the human receiver’s grasp later. As illustrated in Fig. 2, given RGB-D observations of an object on the table, the robot predicts a set of grasp pose candidates together with a human contact map on the object. A final score is computed for each robot grasp pose candidate by combining the confidence score of the grasp and percentage of occluded human contact points. The grasp with the highest score is chosen and executed.
@@ III-A2 Robot Grasp Prediction
Contact-GraspNet is a PointNet++ based U-shaped model that takes in a partial point cloud observation of a scene, and for each point $i$, predicts whether it is contacted by the robot gripper during grasping with a confidence score $S(i)$. For each point $i$, the model predicts the 3-DoF grasp orientation and grasp width $w\in R$ of a parallel-yaw gripper. This 4-DOF grasp representation can then be translated to a 6-DoF robot gripper pose $g$ for each contact point $i$. Following [37], we select grasps with confidence $S(g)\geq 0.23$ as robot grasp candidates. We use the predicted confidence score for each grasp $S(g)$ in the re-ranking phase.
@@ III-A3 Robot Grasp Re-ranking
Clustering Human Contacts: For some objects, like the binoculars, people typically use them with both hands. However, during handovers, a person will likely receive these objects with one hand. Moreover, avoiding all possible contact points might leave few valid robot grasp candidates after re-ranking, leading to grasp failures. To resolve this, we cluster predicted human contact points using DBSCAN[14] based on point spatial density, with no assumptions on the number of clusters. We then choose the largest human contact cluster to compute human contact occlusion and handover orientation in §III-C.
Grasp Re-Ranking: For each grasp $g$, we compute the percentage of human contact points that are occluded by the robot gripper, denoted as $O(g)$. For each human contact point, we ray trace from the contact point along it’s surface normal and check if the ray collides with the robot gripper at grasp pose $g$. If there’s a collision, then we consider the human contact point to be blocked by the gripper at pose $g$. Specifically, $O(g)$ is computed as
$O(g)=\frac{\Big{|}\big{\{}i\in C_{pred}\mid i\textrm{ blocked by }g\big{\}}\Big{|}}{\big{|}C_{pred}\big{|}},$ (1)
@@
Lastly, We re-rank the robot grasps to account for the number of human contacts they occlude. We compute a final contact score $C(g)$ for each grasp, which combines the grasping contact confidence $S(g)$ and the occluded human contacts $O(g)$:
$C(g)=\lambda S(g)-(1-\lambda)O(g).$ (2)
$\lambda$ controls the weight between grasp confidence and human contact occlusion. We use $\lambda=0.5$ in our experiments. Finally, the robot executes the grasp with the highest $C(g)$.
@@ III-B Handover Position
During the delivery phase, the robot should hand the object to a point in front of the human that is both reachable and comfortable for the human arm, with respect to the human’s height and pose. For instance, the handover position for a human that is standing should differ from a human that is sitting down. Following previous works [22, 36, 26], ContactHandover computes the point of handover by minimizing human arm joint torques and joint displacement. In addition, we enforce the handover position to be below the human’s shoulder and above the waist.
@@
Finally, we define the total cost of a candidate point as
@@
where $\alpha$ controls the weight of the two cost functions.
@@ III-C Handover Orientation
In the delivery stage, the robot must present the object in an orientation that minimizes the total distances from contact points to the receiver’s eyes. Specifically, we uniformly sample in the spherical space with a 45 degree granularity and select the object orientations that are kinematically feasible [1] for the robot arm at the computed handover point. For each candidate orientations, ContactHandover computes the total distance between predicted contact points and the human eyes. We estimate the human eye to be located in a position that is exactly 1/2 the height of the human head. The orientation that yields minimum distance is selected.
Note that as in §III-A3, we compute the distance for the largest human contact cluster. For objects with bimodal human contacts (e.g. binoculars), we found that orienting one cluster of human contact towards the human results in more natural handovers, since human tends to receive these objects with one hand instead of two hands. We show the qualitative result for the clustering in §V and Figure 5.
@@ IV-A Handover Metrics
We evaluate our system on 27 daily objects from the ContactDB dataset, which consists of 50 contact maps collected from 50 different users for each object. In this section, we introduce two quantitative metrics based on the contact maps and define success for a robot-human handover.
Metric 1: Human Contact Visibility. To ensure that humans can easily grasp their preferred contact areas upon receiving an object, it is crucial that these areas are visible to the receiver. In particular, the areas where humans prefer to make contact should not be occluded by the object itself, occluded by the robot’s embodiment, or inside the gripper which would be effectively unavailable to the receiver either.
@@
where $CM:i\mapsto c\in\{0,1\}$ indicates if each voxel $i$ on the object surface is touched during a human grasp. $\mathit{P}$ is the set of all voxels in the contact map. $\mathit{V}$ is the set of voxels that are visible from the human’s view without occlusion from the robot embodiment or the object itself, and that are not inside the gripper. We compute visibility by setting a RGB-D camera at the human receiver’s eye position looking at the object; and we consider any contact points that are captured in the image and are not inside the robot gripper to be visible.
@@
where $\mathit{R}=\{i:(d_{1}(i)<l_{\mathbf{arm}})\wedge(d_{2}(i)<d_{2}(\mathbf{gripper}))\}$. $d_{1}$ represents the distance from point $i$ to human shoulder and $l_{\mathbf{arm}}$ is the human’s arm length. $d_{2}(i)$ represents the shortest horizontal distance from point $i$ to the human; $d_{2}(\mathbf{gripper})$ represents the shortest distance between the robot gripper and human. If $d_{2}(i)<d_{2}(\mathbf{gripper})$, then the point $i$ is located between the human and the robot gripper. Higher reachability score indicates more human preferred contact areas are facing towards the human.
Success. For each object, we compute the visibility and reachability score based on the 50 corresponding contact maps and select the median as the final score. We consider a handover to be successful if both visibility and reachability scores exceeds a threshold $k=0.5$.
@@ IV-B Experiment Setup
We set up a revolving RGB-D camera that captures 16 distinct angles around the tabletop to reconstruct the object point cloud and voxel grids. Specifically, we take two set of eight images, one at table height and another at one meter above the table; in both sets, the camera revolves around and looks at the center of the table. We also simulate a RGB-D camera on the human eye level to observe the handover object and compute the visibility metric.
For each object, the robot grasps the object on table, turns around and moves to 1.2 meters in front of the human, and finally handovers the object in the computed pose. We run the experiments for 27 selected objects from the ContactDB “use” dataset under 5 random seeds and report the average.
@@ IV-C Baselines
To evaluate the importance of different components in our approach, we conduct the following ablation studies as shown in Table I:
Ablation 1 (A1): No Grasp Re-ranking (GR). The robot does not predict or consider human contacts when grasping the object. Instead, the robot only predicts and executes the grasp with the highest confidence score.
@@
Ablation 4 (A4). No optimization. The robot executes the grasp with the highest confidence score, moves the robot base to the same location as ContactHandover, while maintaining the same end effector pose after grasping.
Figure 4: Generalize to Unseen Objects. We show ContactHandover’s performance on unseen YCB objects. We show both human contact predictions and handover results on these objects. ContactHandover is able to generalize to (a) objects with unseen shapes (e.g. hammer and knife) and (b) objects with unseen types (e.g. spoon, flat screwdriver, fork and phillips screwdriver). It predicts reasonable human contacts (denoted in red points) around the handles of the objects, picks up and delivers the objects to human with respect to the predicted human contacts.
@@
Figure 5: Clustering bimodal human contacts. For objects with bimodal human contact distributions, ContactHandover clusters the contact points and only optimizes one cluster. In (a) the robot orients one side of the binoculars towards the human. Without clustering, as in (b), both sides are pointed to the human, leaving neither cluster close to the human.